When a clock hand measuring 15.6 cm in length moves through 40 minutes, how far oes the tip move? Give your answer correct to 2 decimal places.
Answer by checkley77(12569)

2*pi*15.6=2*3.14*15.6=97.97 cm for the circumferance.
40/60*97.97=2/3*97.97-65.31 cm is the answer for the distance.